OpenStudy (anonymous):

Find the equation of a line that has a slope of -2 and a y-intercept at (0,9)

OpenStudy (hoblos):

y= -2x +9

OpenStudy (alfie):

\[y = mx +q\] \[y - y_0 = m(x-x_0) \rightarrow y -9 = -2(x) \rightarrow y = -2x+9\]